OpenStudy (anonymous):

A bullet moving at 500 m/s has a kinetic energy of 2500 J and a potential energy of 0.50 at a certain moment. What is the bullet's mass and how high above the ground is it at the moment?

OpenStudy (anonymous):

Hello. First, let's start with what we have. v = 500 m/s k = 2500 J U = 0.50 J \[k = \frac{ 1 }{ 2 }m v^{2}\] \[U = mgh\] To get the mass (m), just use the first equation, since you already have k and v. To get it's height above the surface (h), just use the second equation, since you already have U and m. g is the earth gravitational constant 9.8 m/s^2.

OpenStudy (anonymous):

Ok. Thanks!

OpenStudy (anonymous):

the formula for finding the mass would be m=2k/v^2 ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

and h=U/mg ?

OpenStudy (anonymous):

Yes.